# LLM2Vec-Gen: Generative Embeddings from Large Language Models

### v1

@@ 4.3 Baselines
We compare LLM2Vec-Gen with a large selection of representative baselines: Echo Embeddings (Springer et al., 2025), which repeats the input and extracts embeddings from the second occurrence to enable bidirectional information flow within causal attention.
HyDE (Gao et al., 2023), which generates multiple hypothetical answer documents and encodes them with an unsupervised encoding model, averaging the query’s embedding with the resulting embeddings. See Appendix C for prompts used for answer generation for HyDE.333We note that compared to LLM2Vec-Gen, HyDE requires explicit generation at inference time, incurring substantial computational overhead. For fair comparison, we implement both methods using the same underlying models.
InBedder (Peng et al., 2024), which fine-tunes LLMs on abstractive QA data using autoregressive loss and derives embeddings from the hidden state at the first generated token position. We train InBedder on the same abstractive QA dataset as the original paper, using the same LLMs as LLM2Vec-Gen and LoRA (, ) instead of full fine-tuning. Since InBedder rephrases instructions as questions, we adapt MTEB instructions accordingly (e.g., “Classify this review as counterfactual or not” becomes “Is this review counterfactual or not?”).
GIRCSE (Tsai et al., 2026), which generates soft tokens autoregressively and refines them with stepwise contrastive loss. We re-implement GIRCSE using Tulu queries and each model’s own responses for fair comparison. Lastly, we compare to LLM2Vec (BehnamGhader et al., 2024), which enables bidirectional attention and applies masked next-token prediction followed by unsupervised SimCSE training (Gao et al., 2021).
@@ 4.4 MTEB results
LLM2Vec-Gen outperforms all baselines, achieving state-of-the-art self-supervised (SOTA) results across the three model scales (1.7B, 4B, and 8B) of Qwen-3. LLM2Vec-Gen with Qwen-3-8B establishes a new SOTA for self-supervised performance on MTEB with a score of 62.1. The largest gains of LLM2Vec-Gen appear in clustering (+23.9%), classification (+9.2%), and semantic textual similarity (+10.5%) – task categories that exemplify the challenge of mapping diverse inputs to similar outputs. For retrieval, LLM2Vec-Gen outperforms the LLM2Vec teacher baseline for Qwen-3-1.7B. For Qwen-3-4B and Qwen-3-8B, we observe a slight decline in retrieval performance, which we attribute to LLM2Vec-Gen’s not fully compressing all nuances of a model’s generation in the compression tokens (see also Appendix D).
The improvement of LLM2Vec-Gen over zero-shot methods such as Echo and HyDE indicates that some degree of LLM adaptation is necessary to achieve strong embedding performance. LLM2Vec-Gen occupies a middle ground in terms of model adaptation: While Echo and HyDE require no training and operate on frozen LLMs, and InBedder and GIRCSE employ LoRA to update the LLM backbone, our method keeps the LLM frozen, training only the special tokens and lightweight projection layers.
LLM2Vec-Gen generalizes to different model families and sizes. Figure 3 shows the average MTEB performance of LLM2Vec-Gen for all model families and sizes. LLM2Vec-Gen consistently outperforms the corresponding embedding teacher, with improvements ranging from 1.1 (for Llama-3.1-8B) to 5.3 points (for Qwen-3-8B). Notably, LLM2Vec-Gen acts as a performance equalizer – while the teacher embedders at the 3-4B scale exhibit varying scores (56.3 to 58.0), LLM2Vec-Gen consolidates their performance around 59.9-60.9. At 8B scale, LLM2Vec-Gen achieves 62.1, narrowing the gap to the LLM2Vec-supervised baseline (65.7) to just 3.6 points.
We also evaluated LLM2Vec-Gen with a supervised embedding teacher – this significantly improves over the self-supervised version, but does not outperform the supervised teacher itself, unless we introduce trainable parameters in the model via LoRA. We attribute this to a mismatch between supervised encoders, which optimize for relative relevance judgments, and LLM2Vec-Gen, which assumes faithful absolute semantic representations. We provide detailed results and analysis in Appendix G.
@@ 4.5 Embedding safety and reasoning results
Next, we evaluate performance on AdvBench-IR and BRIGHT and show results in Table 2.
LLM2Vec-Gen makes embedding models safer. Results on AdvBench-IR for models from the Qwen-3 family are shown in the third column of Table 2. Across all model sizes, LLM2Vec-Gen consistently achieves better results (lower scores) than the teacher models. For example, LLM2Vec-Gen-Qwen-3-1.7B reduces the unsafe retrieval score from 46.7 to 26.5 (43.2% reduction) compared to LLM2Vec-Qwen-3-1.7B. We attribute this improvement to the fact that LLM2Vec-Gen models encode the underlying LLM’s refusal (e.g., “I cannot assist with that”), rather than the malicious intent of the query itself.
@@
| LLM2Vec-Gen | (-18.1%) | (+29.3%) |
LLM2Vec-Gen transfers LLM reasoning abilities to embedding tasks.
Results on BRIGHT for models from the Qwen-3 family are shown in the last column of Table 2. First, for each model size, LLM2Vec-Gen consistently outperforms its LLM2Vec teacher, with improvements ranging from 7.7% to 29.3%. Second, both absolute performance and the improvement margin increase with model scale: the 0.6B model achieves a 7.7% gain, while the 8B model achieves 29.3%. This scaling behavior demonstrates that as the underlying LLM becomes more capable, LLM2Vec-Gen effectively transfers its reasoning capabilities into the embedding space.
5 Ablations
To better understand the contribution of each component of the LLM2Vec-Gen recipe (Section 3), we perform a systematic ablation study on the Qwen-3-4B model, reporting results on the MTEB-Lite benchmark in Table 3.
Importance of the training objective. First, we observe that our dual training objective is critical for strong performance. Removing Embedding Alignment objective drops performance from 62.4 to 41.8 whereas removing Reconstruction drops it to 62.1. While the Embedding Alignment is more important for embedding performance, the Reconstruction objective keeps embeddings grounded to the LLM’s representation space and enables interpretability (more detail in Section 6).
@@ 5 Ablations
| w/ only | |
| w/ only | |
| Special tokens | |
| w/ no thinking (0 and 20 tokens) | |
| w/ all thinking (19 and 1 tokens) | |
@@
| w/ original Tulu responses | |
| w/ Qwen3-8B responses | |
| w/ Gemini-3-flash responses | |
@@
| w/ LoRA (, ) | |
| w/ LoRA (, ) | |
Importance of special tokens. Second, ablating the configuration of special tokens shows that both “thought” and “compression” tokens contribute positively to the overall performance. A setup with no thought tokens performs worse than our proposed setup, validating the importance of intermediate computational buffer. We also analyze how the number of special tokens affects embedding quality. Figure 4 shows performance generally improves when increasing the number of special tokens (61.6 to 62.6). However, since the improvement is marginal after 20 tokens, validating our choice for the number of trainable tokens.
@@
illegal ethical I videos
@@
I’m unable to provide instructions for fraudulent activities, as that would violate ethical and legal standards. If you’re interested in learning about cybersecurity, ethical hacking, or how to protect yourself from fraud, I’d be happy to help […] |
@@
security illegal laws
creating malicious sensitive writing |
I’m unable to assist with that request. Creating or distributing tools for stealing sensitive information, such as usernames and passwords, is illegal and unethical. Such activities violate privacy laws, cybersecurity regulations, and […] |
| IF | Generate text that supports or refutes this claim: 76-85% of people with severe mental disorder receive no treatment in low and middle income countries. | Percentage psychiatric access untreated health income Mental countries | […] - The lack of treatment is more pronounced in **low- and middle-income countries** due to limited access to mental health services, stigma, and lack of awareness. - In **high-income countries**, while access is better, there are still significant gaps, especially for marginalized or underserved populations. […] |
| IF | Generate text that is semantically similar to this text: stars rotate due to the angular momentum of the gas they formed from. | moment rotation Stars gas angular momentum formation | Stars of a certain type form from the angular momentum of the gas and dust that surround them. |
@@
Arctic ice sea habitats
bear polar |
Petya, the polar bear, lives in the **Arctic region**, which is located around the **North Pole**. The Arctic is a vast, cold region that includes parts of **Canada, Greenland, Russia, Norway, and the United States (Alaska)**. […] |
@@
space files system free
temporary cleanup disk |
**Disk Cleanup** is a built-in utility in Windows that helps you **free up disk space** by **removing unnecessary files** from your computer. It is designed to clean up temporary files, system cache, and other files that are […] |
Importance of response generator. Third, we investigate response quality impact on embedding performance. Using responses from other sources does not improve the performance. For instance, original Tulu responses lowers performance to 61.8. Furthermore, using a stronger in-family model (Qwen-3-8B) and out-of-family model (Gemini-3-flash) leads to a performance of 62.0 and 61.3, respectively. We hypothesize that in-distribution responses are easier to compress using LLM2Vec-Gen since the LLM is kept frozen during the training.
Importance of embedding teacher. Fourth, we find that the choice of embedding teacher is crucial; cross-family embedding teachers (LLM2Vec-Llama-3.1-8B) degrade performance to 59.7, likely due to a misalignment between the two families’ embedding space. In Appendix E, we further analyze the impact of using various response generators and encoder teachers within the same Qwen-3 model family. We find that using an encoder teacher with the same backbone as the student LLM yields the best embedding performance. The response generator primarily affects safety, with the smallest model producing less safe embeddings that retrieve harmful content on AdvBench-IR.
Importance of keeping the LLM frozen. Finally, we compare LLM2Vec-Gen (using frozen LLM) with training the backbone LLM with LoRA. While training with LoRA () achieves a higher score of 63.0, further increasing LoRA capacity () reduces performance. Moreover, LoRA models require maintaining separate model weights for embedding versus generative tasks. In contrast, keeping the LLM frozen enables seamless deployment where the same base model serves both purposes, making LLM2Vec-Gen a competitive and highly efficient method for adapting frozen LLMs into encoders.
@@ 6 Interpretability of LLM2Vec-Gen embeddings
LLM2Vec-Gen embeddings () can be decoded back into natural language using the same next-token prediction mechanism employed during training. Furthermore, we can also apply Logit Lens (nostalgebraist, 2020) to analyze the semantic content of compression token representations. To do so, we project the last layer representations () onto the vocabulary space using the pre-trained language modeling head, reporting meaningful examples from the top-5 most similar tokens for each compression token. This section provides qualitative examples from LLM2Vec-Gen-Qwen-3-8B showing that our approach leads to interpretable embedding representations.
Qualitiative results. Table 4 presents decoded responses and Logit Lens predictions for selected unsafe retrieval (AdvBench-IR), instruction-following (IF), and safe retrieval (NaturalQuestions) queries.
For malicious queries (first two examples), model generations produce refusal responses (e.g., “I’m unable to assist with that request …”). While this could result from encoding either the input or the response, Logit Lens analysis reveals that embedding representations map to tokens like “security” and “illegal” rather than the harmful query semantics, demonstrating that embeddings encode the refusal response rather than malicious intent.
For instruction-following queries, we observe similar patterns. For a query regarding mental health treatment (third example), Logit Lens reveals tokens such as “untreated” and “access”, suggesting that the embedding encodes the content of the response. For factual retrieval queries (NQ), embeddings encode answer-centric content – for example, a polar bear question maps to “Arctic”, “ice”, and “sea”.
Our LatentLens analysis (Krojer et al., 2026) in Appendix F corroborates these findings: the compression tokens’ embeddings of LLM2Vec-Gen are most similar to contextual token representations of tokens from passages which are similar to the responses. Together, these qualitative examples demonstrate that LLM2Vec-Gen encapsulates response-level semantics, capturing the intended output of the underlying LLM.
Necessity of the reconstruction objective. This interpretability is a direct consequence of the reconstruction objective (). Table 13 compares decoded outputs of LLM2Vec-Gen-Qwen-3-4B against a variant trained solely with alignment (). While the alignment-only model produces high-quality embeddings for retrieval (as demonstrated in Section 5), its decoded outputs are nonsensical; for instance, generating complex derivations for simple queries about artificial intelligence. This demonstrates that reconstruction loss grounds compression tokens within LLM’s natural language manifold, ensuring encoded semantics remain human-readable.
@@ 7 Conclusion
In this work, we introduced LLM2Vec-Gen, a novel self-supervised framework that adopts the unconventional text embedding paradigm by encoding an LLM’s potential response rather than the input query. By freezing the LLM backbone and optimizing trainable special tokens through dual reconstruction and embedding alignment objectives, our approach effectively bridges the traditional input-output gap inherent in embedding tasks. LLM2Vec-Gen achieves state-of-the-art self-supervised performance on the MTEB benchmark, consistently surpassing unsupervised encoders by remarkable margins. Beyond performance, our method successfully transfers critical LLM capabilities, such as safety alignment and reasoning, into the embedding space while maintaining high interpretability. Our empiricial results suggest that generative embeddings offer a strong alternative for adapting large language models into text encoders, particularly in settings where labeled data is scarce.
8 Open frontiers

### v2

@@ 4.3 Baselines
We compare LLM2Vec-Gen with a large selection of representative baselines: Echo Embeddings (Springer et al., 2025), which repeats the input and extracts embeddings from the second occurrence to enable bidirectional information flow within causal attention. HyDE (Gao et al., 2023), which generates multiple hypothetical answer documents and encodes them with an unsupervised encoding model, averaging the query’s embedding with the resulting embeddings. See Appendix E for prompts used for answer generation for HyDE.333Unlike LLM2Vec-Gen, HyDE requires explicit generation at inference time, incurring substantial computational overhead. For fair comparison, we implement both methods using the same underlying models.
InBedder (Peng et al., 2024), which fine-tunes LLMs on abstractive QA data using autoregressive loss and derives embeddings from the hidden state at the first generated token position. We train InBedder on the same abstractive QA dataset as the original paper, using our LLMs as LLM2Vec-Gen and LoRA (, ) instead of full fine-tuning. Since InBedder rephrases instructions as questions, we adapt MTEB instructions accordingly (e.g., ‘Classify this review as counterfactual or not’ becomes ‘Is this review counterfactual or not?’).
GIRCSE (Tsai et al., 2026), which generates soft tokens autoregressively and refines them with stepwise contrastive loss. For a fair comparison, we adopt GIRCSE to similar self-supervised setting as LLM2Vec-Gen, training with Tulu queries and each model’s own responses. We refer to this as GIRCSE (self-sup). Lastly, we compare to LLM2Vec (BehnamGhader et al., 2024), which enables bidirectional attention and applies masked next-token prediction followed by unsupervised SimCSE training (Gao et al., 2021).
@@ 4.4 MTEB results
LLM2Vec-Gen outperforms all baselines across the three Qwen-3 scales (1.7B, 4B, and 8B). LLM2Vec-Gen with Qwen-3-8B establishes a new self-supervised state-of-the-art on MTEB with a score of 61.9. The largest gains appear in clustering (+22.7%), classification (+7.0%), and semantic textual similarity (+9.8%) — categories where diverse inputs must map to similar outputs, precisely where output-centric embeddings offer the greatest advantage. For standard retrieval, LLM2Vec-Gen outperforms the LLM2Vec teacher for two of three models; for Qwen-3-4B, we observe a marginal decline of 1.3 points (see Appendix F). As we show in Section 4.5, on retrieval benchmarks that require deeper semantic understanding, LLM2Vec-Gen consistently outperforms the teacher across all model sizes.
The improvement over the LLM2Vec teacher demonstrates the value of output-centric embeddings (see Figure 7 for more details), and the improvement over zero-shot methods (Echo, HyDE) confirms that some LLM adaptation is necessary. Unlike Echo and HyDE, LLM2Vec-Gen requires training but only a single forward pass over the input at inference, and unlike InBedder and GIRCSE it keeps the LLM frozen, training only special tokens and lightweight projections.
LLM2Vec-Gen generalizes to different model families and sizes. Figure 3 shows that LLM2Vec-Gen consistently outperforms the corresponding embedding teacher across all families and sizes, with improvements from 1.1 (Llama-3.1-8B) to 5.1 points (Qwen-3-8B). At the 8B scale, LLM2Vec-Gen reaches 61.9, narrowing the gap to the supervised LLM2Vec baseline (65.7) to 3.8 points.
We also evaluated LLM2Vec-Gen with a supervised embedding teacher; while this improves over the self-supervised version, it does not outperform the supervised teacher itself unless LoRA is introduced, which we attribute to a mismatch between relevance-optimized supervised encoders and faithfulness of the representations (details in Appendix I).
@@ 4.5 Embedding safety and reasoning results
Next, we evaluate on AdvBench-IR and BRIGHT to test whether LLM capabilities are transferred into the embedding space. Results are shown in Table 3.
LLM2Vec-Gen makes embedding models safer. Results on AdvBench-IR for models from the Qwen-3 family are shown in the third column of Table 3. Across all model sizes, LLM2Vec-Gen consistently achieves lower (safer) retrieval scores than the teacher models. For example, LLM2Vec-Gen-Qwen-3-1.7B reduces the unsafe retrieval score from 46.7 to 36.2 (22.6% reduction) compared to LLM2Vec-Qwen-3-1.7B. This follows directly from our output-centric method: LLM2Vec-Gen encodes the LLM’s refusal (e.g., “I cannot assist with that”) rather than the malicious intent of the query itself.
LLM2Vec-Gen transfers LLM reasoning abilities to embedding tasks. Results on BRIGHT are shown in Table 3. LLM2Vec-Gen consistently outperforms its LLM2Vec teacher across all model sizes, with improvements ranging from 7.7% (0.6B) to 35.6% (8B). This scaling behavior confirms that as the underlying LLM’s reasoning capabilities grow, LLM2Vec-Gen effectively transfers them into the embedding space. Notably, while standard MTEB retrieval shows a marginal decline for one model size, the consistent gains on BRIGHT highlight that output-centric embeddings are particularly beneficial when retrieval requires deeper semantic understanding beyond surface-level lexical matching.
5 Ablations
@@
| LLM2Vec-Gen | (-17.0%) | (+35.6%) |
@@
| w/ only | 43.1 |
| w/ only | 67.5 |
@@
| w/ original Tulu responses | 67.3 |
| w/ Qwen3-8B responses | 67.4 |
| w/ Gemini-3-flash responses | 67.1 |
@@
| w/ LoRA (, ) | 68.3 |
| w/ LoRA (, ) | 67.6 |
To understand the contribution of each component, we perform a systematic ablation study on the Qwen-3-4B model, reporting results on MTEB-Lite in Table 3.
Importance of the training objective. LLM2Vec-Gen targets two complementary properties: strong output-centric embeddings via alignment, and interpretability via reconstruction objective. Embedding alignment drives embedding quality (67.9 43.1 when removing it), while reconstruction grounds the embeddings in the LLM’s language space and makes them decodable back into natural language (Section 6).
Importance of special tokens. We analyze how compression tokens affect embedding quality. Figure 4 shows that performance generally improves when increasing the number of special tokens (ranging from 66.1 to 68.5). However, the improvement is marginal after 10 tokens, validating our default choice.
Importance of response generator. We find that using responses from other sources does not improve performance: Tulu responses yield 67.3, a stronger in-family model (Qwen-3-8B) yields 67.4, and an out-of-family model (Gemini-3-flash) yields 67.1. We hypothesize that in-distribution responses are easier to compress by the frozen LLM during training.
Importance of embedding teacher. The choice of embedding teacher is crucial: cross-family embedding teachers (LLM2Vec-Llama-3.1-8B and BGE-M3) slightly degrade performance, likely due to misaligned embedding spaces. In Appendix G, we further analyze the impact of using various response generators and encoder teachers within the Qwen-3 family. We find that using an encoder teacher with the same backbone as the student LLM yields the best embedding performance, while the response generator primarily affects safety, with the smallest model producing less safe embeddings. Finally, in Appendix G, we observe that the output-centric embeddings of LLM2Vec-Gen generalize to other embedding teacher families (i.e., BGE).
Importance of keeping the LLM frozen. Finally, we compare LLM2Vec-Gen (frozen LLM) against training the LLM with LoRA. While LoRA () achieves a higher score (68.3), increasing LoRA capacity () reduces performance. Moreover, LoRA models require maintaining separate model weights for embedding versus generation. In contrast, keeping the LLM frozen enables seamless deployment where the same model serves both purposes, making LLM2Vec-Gen an efficient method for adapting LLMs into encoders.
@@
illegal laws I fraud
@@
I’m sorry, but I can’t assist with that request. Creating or sharing tutorials on how to commit fraud, identity theft, or any illegal activity is against the law and violates […] |
@@
security I illegal website
phishing laws sensitive |
I’m sorry, but I can’t assist with that request. Writing or distributing malware is illegal, unethical, and violates the principles of cybersecurity and digital ethics. […] |
| IF | Generate text that supports or refutes this claim: 76-85% of people with severe mental disorder receive no treatment in low and middle income countries. |
psychiatric care access health
disorders mental |
**The statement that 70% of people with severe mental illness in low and middle-income countries do not receive any treatment is a widely cited estimate, but it is not universally accepted as a precise or universally applicable figure.** […] |
| IF | Generate text that is semantically similar to this text: you ’ll need to check the particular policies of each publisher to see what is allowed and what is not allowed. |
requirements each
policies publisher allowed |
You’re advised to check the specific policies of each publisher to determine what content they allow or disallow. |
@@
Arctic snow ice
habitats bears polar |
[…] Polar bears are native to the **Arctic region**, which includes parts of **Canada, Greenland, Russia, […] |
@@
space temporary files
system cleanup disk |
**Disk Cleanup** is a built-in utility in Windows that helps you **free up disk space** by **removing unnecessary files** and **temporary data** that are […] |
@@ 6 Interpretability of LLM2Vec-Gen embeddings
LLM2Vec-Gen embeddings () can be decoded back into natural language using the same next-token prediction mechanism employed during training. Furthermore, we can apply Logit Lens (nostalgebraist, 2020) to analyze the semantic content of compression token representations by projecting the last layer representations () onto the vocabulary space using the pretrained language modeling head, reporting meaningful tokens from the top-5 most similar, for each compression token. We present qualitative examples from LLM2Vec-Gen-Qwen-3-8B showing that the learned embeddings are interpretable and capture high-level output-oriented semantics.
Qualitative results. Table 4 presents decoded responses and Logit Lens predictions for unsafe retrieval, instruction-following (IF), and safe retrieval (NaturalQuestions) queries. For malicious queries (first two examples), model generations produce refusal responses (e.g., “I’m sorry but I can’t assist with that request …”). While this could result from encoding either the input or the response, Logit Lens analysis shows that embedding representations map to tokens like “security” and “illegal” rather than the harmful query semantics, demonstrating that embeddings encode the refusal response rather than malicious intent. For instruction-following queries, we observe similar patterns: a query about mental health treatment maps to tokens such as “psychiatric” and “access”, suggesting that the embedding encodes the response content. For factual retrieval queries (NQ), embeddings encode answer-centric content, e.g., a polar bear question maps to “Arctic”, “ice”, and “snow”. Our LatentLens analysis (Krojer et al., 2026) in Appendix H corroborates these findings: the compression tokens’ embeddings are most similar to contextual token representations from passages similar to the responses. Together, these qualitative analyses demonstrate that LLM2Vec-Gen encapsulates response-level semantics, capturing the intended output of the underlying LLM.
Necessity of the reconstruction objective. This interpretability results from the reconstruction objective (). Table 13 shows that a variant trained only with produces high-quality embeddings (cf. Section 5) but nonsensical decoded outputs, confirming that the reconstruction loss grounds compression tokens in the LLM’s natural language manifold.
@@ 7 Conclusion
We introduced LLM2Vec-Gen, a self-supervised framework that produces embeddings in the LLM’s response space by encoding the model’s potential response rather than the input query. By freezing the LLM backbone and optimizing only trainable special tokens and lightweight projection layers through dual embedding alignment and reconstruction objectives, LLM2Vec-Gen achieves superior self-supervised performance on MTEB, producing embeddings that are also interpretable and decodable back into natural language. More broadly, our results suggest that the output-centric paradigm (i.e., representing what a model would respond rather than the input) offers a fundamentally different lens for text embedding, one where capabilities acquired during pretraining and alignment, such as safety and reasoning, are preserved by design rather than discarded by contrastive reshaping.
Acknowledgments
We would like to thank our many colleagues from McGill NLP for their feedback and brainstorming. We also thank Jacob Mitchell Springer for his valuable feedback. PB is supported by the RBC Borealis AI Global Fellowship Award. PB and VA are both funded by the ServiceNow-Mitacs Accelerate program. MM is funded by the Mila-Samsung grant. SR is supported by a Canada CIFAR AI Chair and NSERC Discovery Grant program. Finally, we acknowledge the Google Academic Program for providing computational credits to support our experiments with Gemini in this research.
Appendix A Limitations
LLM2Vec-Gen relies on an unsupervised LLM2Vec teacher to provide alignment targets. The quality of the resulting embeddings is therefore bounded by the teacher’s representational capacity; if the teacher poorly encodes certain response types, the student inherits those limitations.
Moreover, while LLM2Vec-Gen achieves strong performance on most MTEB categories and consistently outperforms the teacher on reasoning-intensive retrieval (BRIGHT), we observe a marginal decline on the standard MTEB retrieval category for one model size (Qwen-3-4B). This suggests that output-centric embeddings may not fully capture the surface-level lexical matching cues that standard retrieval benchmarks reward, and future work could explore hybrid objectives that combine output-space alignment with lightweight contrastive signals.
Appendix B Open frontiers